Abstract

Suppose c ≥ 2 and d ≥ 2 are integers, and let S be the set of integers ⌊c^{j}/d^{k}⌋, where j and k range over the nonnegative integers. Assume that c and d are multiplicatively independent; that is, if p and q are integers for which c^{p} = d^{q}, then p = q = 0. The numbers in S form interspersions in various ways. Related fractal sequences and permutations of the set of nonnegative integers are also discussed.